Consider the set V = R 2 . For (x1, y1),(x2, y2) ∈ R 2 and c ∈ R, define the following operations:

I. (x1, y1) + (x2, y2) = (0, y1 + y2)

II. c(x1, y1) = (0, cy1)

Is the subset a vector of R2. If not prove the axioms that makes it false.Also prove those axioms that are true.


1
Expert's answer
2022-03-19T02:40:21-0400

"V=R^2\\
I. \forall \vec{a}=(x_1,y_1)\in V, \vec{b}=(x_2,y_2)\in V\\
\vec{a}+\vec{b}=(x_1,y_1)+(x_2,y_2)=(0,y_1+y_2)\in V\\
II. \forall c\in R, \forall \vec{a}=(x_1,y_1)\in V\\
c\cdot\vec{a}=c\cdot(x_1,y_1)=(0,cy_1)\in V"

Axioms

"1) \vec{a}+\vec{b}= \vec{b}+\vec{a}\\
 \vec{a}+\vec{b}=(x_1,y_1)+(x_2,y_2)=(0,y_1+y_2)\\
 \vec{b}+\vec{a}=(x_2,y_2)+(x_1,y_1)=(0, y_2+y_1)=\\
=(0,y_1+y_2)= \vec{a}+\vec{b}\\
true"

"2. \forall \vec{d}=(x_3,y_3)\in V\\
(\vec{a}+\vec{b})+\vec{d}=\vec{a}+(\vec{b}+\vec{d})\\
\vec{a}+(\vec{b}+\vec{d})=(x_1,y_1)+(0, y_2+y_3)=\\
=(0, y_1+y_2+y_3)\\
(\vec{a}+\vec{b})+\vec{d}=(0, y_1+y_2)+(x_3,y_3)=\\
=(0, y_1+y_2+y_3)=\vec{a}+(\vec{b}+\vec{c})\\
true"

"3. \exist\vec{0}=(x_2,y_2)\in V, \forall\vec{a}=(x_1,y_1)\in V\\
\vec{a}+\vec{0}=\vec{a}\\
\vec{a}+\vec{0}=(x_1,y_1)+(x_2,y_2)=(0, y_1+y_2)\neq\vec{a}\\
\not\exist\vec{0}\\
false"

"4. \forall\vec{a}=(x_1,y_1)\in V\exist\vec{b}=(x_2,y_2)\in V\\
\vec{a}+\vec{b}=\vec{0}\\
\not\exist\vec{0}\\
false"

"5. \forall c, k\in R, \forall \vec{a}=(x_1,y_1)\in V\\
(c+d)\cdot\vec{a}=c\cdot\vec{a}+d\cdot\vec{a}\\
(c+d)\cdot\vec{a}=(c+d)\cdot(x_1,y_1)=(0,(c+d)y_1)\\
c\cdot\vec{a}=c\cdot(x_1,y_1)=(0,cy_1)\\
d\cdot\vec{a}=d\cdot(x_1,y_1)=(0,dy_1)\\
c\cdot\vec{a}+d\cdot\vec{a}=(0,cy_1)+(0,dy_1)=(0,cy_1+dy_1)=\\
=(c+d)\cdot\vec{a}\\true"

"6. \forall c\in R, \forall \vec{a}=(x_1,y_1)\in V,\\
\vec{b}=(x_2,y_2)\in V\\
c\cdot(\vec{a}+\vec{b})=c\cdot\vec{a}+c\cdot\vec{b}\\
c\cdot(\vec{a}+\vec{b})=c\cdot(0,y_1+y_2)=(0,c(y_1+y_2))\\
c\cdot\vec{a}=(0,cy_1)\\
c\cdot\vec{b}=(0,cy_2)\\
c\cdot\vec{a}+c\cdot\vec{b}=(0,cy_1)+(0,cy_2)=(0,cy_1+cy_2)=\\
=c\cdot(\vec{a}+\vec{b})\\true"

"7. \forall c, k\in R, \forall \vec{a}=(x_1,y_1)\in V\\
(c\cdot d)\cdot\vec{a}=c\cdot(d\cdot\vec{a})\\
(c\cdot d)\cdot\vec{a}=(c\cdot d)\cdot(x_1,y_1)=(0,c\cdot dy_1)\\
d\cdot\vec{a}=d\cdot(x_1,y_1)=(0,dy_1)\\
c\cdot(d\cdot\vec{a})=c\cdot(0,dy_1)=(0,cdy_1)=
(c\cdot d)\cdot\vec{a}\\true"

"8. \exist 1\in R, \forall\vec{a}=(x_1,y_1)\in V\\
1\cdot\vec{a}=\vec{a}\\
1\cdot\vec{a}=1\cdot(x_1,y_1)=(0,1\cdot y_1)=(0,y_1)\neq\vec{a}\\
false"

"V" is not the subset a vector of R2